# Ryoga Matsuo2,3, Stefan Uhlich3, Arun Venkitaraman1, Andrea Bonetti1, Chia-Yu Hsieh1, Ali Momeni1,2 Lukas Mauch3, Augusto Capone3, Eisaku Ohbuchi4, Lorenzo Servadei1 1SonyAI, Switzerland2EPFL, Switzerland3Sony Semiconductor Solutions Europe, Germany4Sony Semiconductor Solutions, Japan

### v1

@@ IV-D Performance metrics
We use the following three metrics to evaluate the samples generated by the different models: (a) BLEU that measures an N-gram similarity, (b) CSR that measures syntatic correctness, and (c) MSSIM that measures similarity with human generated design. We describe these metrics next.
IV-D1 Text-based metric
The bilingual evaluation understudy (BLEU) compares N-grams of the generated text and the reference translation and counts the number of matches independently of the positions [25]. With this measure, we compare the N-gram similarity between the reference and generated text-based schematics. While BLEU is not perfectly suited for our tasks due to the syntactical nature of .asc and files, it can indicate the extent to which the pre-trained models already understand their language syntax. Furthermore, we selected the standard BLEU as our metric over CodeBLEU [26] specialized to capture syntactic and semantic features of codes since this measure does not support .asc and formats. We compare the performance of the models by computing $\overline{\text{BLEU}}$, that is the average BLEU score across all test samples.
IV-D2 Compilation-based metric
The compilation success rate (CSR) measures the syntactical correctness of the generated .asc and scripts and is defined as
$\mbox{CSR}=\displaystyle\frac{N_{\text{compilable}}}{N_{\text{ref}}}$
where $N_{\text{compilable}}$ is the total number of compilable generated codes in the test set and $N_{\text{ref}}$ is the total number of reference samples.
IV-D3 Image-based metric
The structural similarity index measure (SSIM) is a metric that computes the structural similarity of two image signals based on the statistical comparisons in luminance, contrast, and structure [27]. For a given pair of image patches $x$ and $y$, SSIM is defined as
@@ IV-D3 Image-based metric
where $x_{i}$ and $y_{i}$ are the image signals in the $i$-th local window and $M$ is the number of local windows applied on the images. We use this metric to derive the similarities between the reference schematics in the test set and the schematics generated by LLMs from netlists. To obtain the images of generated and reference schematics, we load each .asc file with LTSpice and take a screenshot. These images are subsequently cropped to eliminate the artifacts from the graphical user interface. For the case of codes, the .tex files are converted into PDF files using pdflatex. These are then converted into images with pdf2image [^3] https://github.com/Belval/pdf2image. For cases where the generated codes do not successfully compile into an image, we set the MSSIM to be equal to 0. We then compare the performance of the models using $\overline{\mbox{MSSIM}}$, that is the MSSIM score averaged over all test samples.
@@ V Results & Discussion
In order to select the starting model for $Schemato$, we evaluated the performance of Llama-3.1-8B raw models across various configurations. Table I shows the performance of the different models for the netlist-to-.asc conversion task. We note that the $\overline{\text{MSSIM}}$ values are normalized by the maximum $\overline{\text{MSSIM}}$ value of 0.0228 achieved by $Schemato$ fine-tuned on prompt 4. As can be seen, the performance did not vary significantly across different settings. Notably, moving from zero-shot to one-shot prompting had negligible impact on the performance for this task. In contrast, Table II highlights a key observation for the netlist-to- conversion task. Again, we note that the $\overline{\text{MSSIM}}$ values are normalized by the maximum $\overline{\text{MSSIM}}$ value of 0.0217 achieved by $Schemato$ fine-tuned on prompt 4 for the corresponding task. Here, Llama-3.1-8B instruct with prompt 2 achieved a CSR of 25% for generation, while all other raw configurations failed to produce any compilable outputs. Based on these observations, we selected Llama-3.1-8B instruct with prompt 2 as the starting configuration. Additionally, we fine-tune Llama-3.1-8B instruct with prompt 4 to investigate the effect of one-shot prompting.
During fine-tuning, training and validation losses (Fig. 3) revealed overfitting for the netlist-to-.asc conversion after epoch 2 (since the validation loss increased with training epochs), while netlist-to- validation loss stabilized after several epochs. Since the cross entropy loss does not directly help quantify the correctness of the code and the generated schematics, we use CSR, and $\overline{\text{MSSIM}}$ metrics discussed in Section IV-D to evaluate the model performance during training epochs as shown in Fig. 4. We then choose the model weights that gave the best pair of CSR and $\overline{\text{MSSIM}}$ scores for further evaluation.
@@
(a) instruct, prompt 2
(b) instruct, prompt 4
Figure 3: Training and validation losses of $Schemato$ fine-tuned on zero-shot and one-shot prompting. Validations losses are measured at each epoch.
(a) instruct, prompt 2
(b) instruct, prompt 4
Figure 4: CSR and $\overline{\text{MSSIM}}$ of $Schemato$ fine-tuned with zero-shot and one-shot prompting for netlist-to- conversion.
For the netlist-to-.asc conversion task, we show the reference and generated schematics for a circuit example of an NPN bipolar transistor circuit from $Schemato$ in Fig. 5 for the different models. Similarly, for the netlist-to- conversion task, we show the schematics obtained for the example of an RC phase shift network in Fig. 6. We note that the raw models of Llama-3.1-8B failed to produce a compilable code for the second example. Although $Schemato$’s schematics differ geometrically from the reference, they maintain identical topological features, demonstrating successful netlist-to-schematic conversion. GPT-4o with prompt 5 produced the correct RC network topology but with overlapping components, whereas $Schemato$ generated a clean schematic with well-spaced components, adhering to human design principles. These examples underscore $Schemato$’s ability to produce interpretable and accurate schematics, even without geometrical guidance, surpassing Llama-3.1-8B and GPT-4o in performance.
Q1 N001 N002 0 0 NPN
@@
with prompt 2
(d) Llama 3.1 base
with prompt 2
(e) GPT-4o
with prompt 4
Figure 5: Netlist & LTSpice schematics of an NPN transistor circuit.
@@
(c) $Schemato$ with prompt 4
(d) GPT-4o with prompt 4
(e) GPT-4o with prompt 5
Figure 6: Netlist & schematics of an RC phase shift network.
TABLE I: Test Performance for netlist-to-.asc conversion. The best and second best metrics are highlighted in red and blue, respectively. $\overline{\text{MSSIM}}$ values are normalized by the maximum of 0.0228.
Model Version Prompt $\mathbf{\overline{BLEU}}$ CSR (%) $\mathbf{\overline{MSSIM}}$ $\mathbf{N_{\text{ref}}}$
Without Fine–tuning: Zero shot
Llama-3.1-8B base 1 0.00 4.08 (2/49) 0.0395 49
Llama-3.1-8B instruct 1 0.00 0.00 (0/49) 0.0000 49
GPT 4o 1 8.40 16.33 (8/49) 0.1140 49
Llama-3.1-8B base 2 0.00 20.41 (10/49) 0.1798 49
Llama-3.1-8B instruct 2 1.51 12.24 (6/49) 0.1754 49
GPT 4o 2 10.87 32.65 (16/49) 0.2237 49
Llama-3.1-8B base 3 0.00 20.41 (10/49) 0.1711 49
Llama-3.1-8B instruct 3 0.00 10.20 (5/49) 0.1535 49
GPT 4o 3 9.97 24.49 (12/49) 0.1798 49
Without Fine–tuning: One-shot
Llama-3.1-8B base 4 0.00 20.41 (10/49) 0.1711 49
Llama-3.1-8B instruct 4 5.08 32.65 (16/49) 0.2158 49
GPT 4o 4 13.65 48.98 (24/49) 0.3772 49
Llama-3.1-8B base 5 0.00 24.49 (12/49) 0.1974 49
Llama-3.1-8B instruct 5 1.18 20.41 (10/49) 0.1272 49
GPT 4o 5 14.46 73.47 (36/49) 0.5833 49
Fine-tuned: epoch 3 & 2, respectively
Schemato instruct 2 12.33 71.43 (35/49) 0.7158 49
Schemato instruct 4 13.43 63.27 (31/49) 1.0000 49
TABLE II: Test Performance for netlist-to- conversion. The best and second best metrics are highlighted in red and blue, respectively. $\overline{\text{MSSIM}}$ values are normalized by the maximum of 0.0217.
Model Version Prompt $\mathbf{\overline{BLEU}}$ CSR (%) $\mathbf{\overline{MSSIM}}$ $\mathbf{N_{\text{ref}}}$
Without Fine–tuning: Zero shot
Llama-3.1-8B base 1 0.16 0.00 (0/36) 0.0000 36
Llama-3.1-8B instruct 1 0.00 0.00 (0/36) 0.0000 36
GPT 4o 1 0.01 0.00 (0/49) 0.0000 49
Llama-3.1-8B base 2 0.00 0.00 (0/36) 0.0000 36
Llama-3.1-8B instruct 2 3.08 25.00 (9/36) 0.2622 36
GPT 4o 2 0.00 0.00 (0/49) 0.0000 49
Llama-3.1-8B base 3 0.93 0.00 (0/36) 0.0000 36
Llama-3.1-8B instruct 3 4.79 0.00 (0/36) 0.0000 36
GPT 4o 3 0.05 0.00 (0/49) 0.0000 49
Without Fine-tuning: One-shot
Llama-3.1-8B base 4 9.08 0.00 (0/32) 0.0000 32
Llama-3.1-8B instruct 4 47.55 0.00 (0/32) 0.0000 32
GPT 4o 4 16.18 24.49 (12/49) 0.3221 49
Llama-3.1-8B base 5 7.55 0.00 (0/32) 0.0000 32
Llama-3.1-8B instruct 5 48.99 0.00 (0/32) 0.0000 32
GPT 4o 5 21.96 26.53 (13/49) 0.1751 49
Fine-tuned: epoch 7 & 9, respectively
Schemato instruct 2 69.90 86.11 (31/36) 0.7272 36
Schemato instruct 4 65.04 93.75 (30/32) 1.0000 32
@@ VI Conclusions & Future work
In this paper, we introduced and evaluated $Schemato$, a fine-tuned LLM designed for converting circuit netlists into interpretable visual schematics. Our experimental results demonstrate that $Schemato$ significantly outperforms state-of-the-art LLMs, including Llama-3.1-8B and GPT-4o, across all the text-based, compilation-based, and image-based metrics in the netlist-to- conversion task. Similarly, for the netlist-to-.asc conversion tasks, $Schemato$ outperforms all the other models in the image-based score ($\overline{\text{MSSIM}}$). These findings underscore the potential of $Schemato$ in reliably generating human-intuitive schematics from netlists, paving the way for enhanced interpretability and utility in ML-based analog circuit synthesis.
Our experiments show that for more complex examples, such as a NAND2 and a transistor-level OpAmp, $Schemato$ struggles to generate accurate schematics. This is expected as our training dataset with 2,712 different circuits lacks structural diversity. This highlights the need for a larger dataset to better capture the nuances of such designs. To address this challenge, future improvements to $Schemato$ will focus on enriching the dataset using additional circuits and advanced data augmentation techniques. For instance, we could combine two schematics into a new one using a technique similar to mixup used in vision [28]. To improve the generalizability of $Schemato$ to large and unseen circuit classes, we plan to investigate the decomposition of larger circuits into smaller, more manageable subcircuits, thereby enhancing the ability of the model to learn hierarchical structures and the interplay between subcircuits. A second limitation of our current work is that the image-based evaluation metric does not consider the circuits’ topological structure and is highly sensitive to absolute locations of circuit components in the image. This resulted in the very small maximum $\overline{\text{MSSIM}}$ values and may lead to an unfair comparison between the reference and generated schematics. A possible direction to address this is to investigate the use of graph representations and graph-based metrics that capture the topological information from the generated schematic. We intend to pursue these directions in the future.

### v2

@@ IV-D Performance metrics
We use the following four metrics to evaluate the samples generated by the different models in order to capture various aspects of schematic generation: (1) graph edit distance (GED) score that measures the similarity of ground truth and generated graphs to compare their connectivity, (2) mean structural similarity index measure (MSSIM) that measures visual similarity with human generated design, (3) compilation success rate (CSR) that measures syntatic correctness, and (4) bilingual evaluation understudy (BLEU) that measures an N-gram similarity. We describe these metrics next.
IV-D1 Graph-based metric
The graph edit distance (GED) measures the minimum number of modifications (inserting/deleting/substituting a node/edge) required to transform one graph to another. The algorithm to compute this metric is often implemented with shortest-path algorithms like the Dijkstra’s algorithm [26]. As this metric is provided as an integer value, we need to normalize it such that it is agnostic to the graph size. We formulate the GED score as
GED $\text{GED score}=1-\frac{\text{GED}}{\text{max}(N_{\text{n, gen}},N_{\text{n, ref}})+\text{max}(N_{\text{e, gen}},N_{\text{e, ref}})}$
where $N_{\text{a, b}}$ is the number of a (n for nodes and e for edges) in b (gen for generated and ref for reference) samples. The computation of GED relies on the function, “graph_edit_distance” from NetworkX [27], which utilizes the depth-first graph edit distance algorithm [28]. The time-out is set to 60 seconds for each sample pair comparison to prevent an excessively long search. The function returns the minimum GED found so far once this limit is reached. To map the generated asc files into the form of graphs, we again use LTSpice to compile them back into netlists. Then, we convert these netlists and the reference netlists provided in the prompt into nxgraph class using NetworkX and compute the GED scores. For cases where the generated asc script does not successfully compile into a netlist, we omit the sample from the GED score calculation. We then compare the performance of the models using GED $\overline{\mbox{GED score}}$, that is the GED score averaged over all compilable test samples.
IV-D2 Image-based metric
The structural similarity index measure (SSIM) is a metric that computes the structural similarity of two images based on the statistical comparisons in luminance, contrast, and structure [29]. For a given pair of image patches $x$ and $y$, SSIM is defined as
@@ IV-D2 Image-based metric
where $x_{i}$ and $y_{i}$ are the image patches in the $i$-th local window and $M$ is the number of local windows applied on the images. We use this metric to derive the similarities between the reference schematics in the test set and the schematics generated by LLMs from netlists. To obtain the images of generated and reference schematics, we load each .asc file with LTSpice and take a screenshot. These images are subsequently cropped to eliminate the artifacts from the graphical user interface. Similarly to the GED scores, we omit the incompilable samples from the calculation. We then compare the performance of the models using $\overline{\mbox{MSSIM}}$, the MSSIM score averaged over all compilable test samples.
IV-D3 Compilation-based metric
The compilation success rate (CSR) measures the syntactical correctness of the generated .asc files and is defined as
$\mbox{CSR}=\displaystyle\frac{N_{\text{compilable}}}{N_{\text{ref}}}$
where $N_{\text{compilable}}$ is the total number of generated schematics in the test set that can be compiled into netlists using LTSpice and $N_{\text{ref}}$ is the total number of reference samples.
IV-D4 Text-based metric
The bilingual evaluation understudy (BLEU) compares N-grams of the generated text and the reference translation and counts the number of matches independently of the positions [30]. With this measure, we compare the N-gram similarity between the reference and generated text-based schematics. While BLEU is not perfectly suited for our tasks due to the syntactical nature of .asc files, it can indicate the extent to which the pre-trained models already understand their file syntax. Furthermore, we selected the standard BLEU as our metric over CodeBLEU [31] specialized to capture syntactic and semantic features of codes since this measure does not support .asc file formats. We compare the performance of the models by computing $\overline{\text{BLEU}}$, that is the average BLEU score across all test samples.
TABLE I: Test performance comparison. $\overline{\text{GED}}$ is equivalent to the average GED score, GED $\overline{\text{GED score}}$. The values on the left and right side of / indicate the raw scores and scores scaled by CSR, respectively for GED $\overline{\text{GED score}}$ and $\overline{\text{MSSIM}}$. The best and second best metrics are highlighted in red and blue respectively.
Model Version Prompt $\mathbf{\overline{GED}}$ $\mathbf{\overline{MSSIM}}$ CSR(%) $\mathbf{\overline{BLEU}}$
Pretrained models
Llama-3.1-8B base 1 - - 0.00 0.00
Llama-3.1-8B instruct 1 - - 0.00 0.00
GPT 4o 1 0.21/0.06 0.05/0.01 27.35 9.24
Llama-3.1-8B base 2 0.26/0.01 0.02/0.00 3.42 6.40
Llama-3.1-8B instruct 2 0.26/0.01 0.02/0.00 3.42 6.40
GPT 4o 2 0.22/0.06 0.04/0.01 27.35 11.97
Llama-3.1-8B base 3 - - 0.00 0.00
Llama-3.1-8B instruct 3 - - 0.00 0.00
GPT 4o 3 0.13/0.00 0.02/0.00 0.85 8.76
Llama-3.1-8B base 4 0.30/0.13 0.06/0.03 41.88 10.51
Llama-3.1-8B instruct 4 0.30/0.13 0.03/0.01 41.88 10.51
GPT 4o 4 0.23/0.12 0.07/0.04 53.85 15.22
Llama-3.1-8B base 5 0.27/0.10 0.06/0.02 38.46 10.55
Llama-3.1-8B instruct 5 0.27/0.10 0.06/0.02 38.46 10.55
GPT 4o 5 0.23/0.15 0.03/0.02 63.25 14.84
Best fine-tuned model
$Schemato$ instruct 3 0.35/0.27 0.22/0.17 76.07 22.17
TABLE II: Ablation study for base model selected for fine-tuning. The values on the left and right side of / indicate the raw scores and scores scaled by CSR, respectively for GED $\overline{\text{GED score}}$ and $\overline{\text{MSSIM}}$. For each prompt, the best performing weights are chosen. The best metrics are highlighted in red.
Model Epoch Version Prompt $\mathbf{\overline{GED}}$ $\mathbf{\overline{MSSIM}}$ CSR(%) $\mathbf{\overline{BLEU}}$
Models fine-tuned with different prompts
Llama-3.1-8B 4 instruct 2 0.39/0.27 0.03/0.02 70.09 20.11
Llama-3.1-8B 5 instruct 4 0.43/0.28 0.04/0.03 64.10 15.26
Llama-3.1-8B 9 instruct 5 0.33/0.26 0.20/0.16 77.78 23.40
Best fine-tuned model
$Schemato$ 6 instruct 3 0.35/0.27 0.22/0.17 76.07 22.17
Figure 3: Training and validation losses. Llama-3.1-8B is fine-tuned with four different prompts: (a) Prompt 2, (b) Prompt 3, (c) Prompt 4, and (d) Prompt 5. Validations losses are computed after every training epoch.
@@ V Results & Discussion
In order to select the starting model for fine-tuning, we evaluated the performance of Llama-3.1-8B raw models across various configurations. Table I shows the performance of the different models on the evaluation using the test set. As can be seen, the performance did not vary across the two versions: base and instruct. Moving from zero-shot (Prompt 1-3) to one-shot prompting (Prompt 4-5) improved the best CSR significantly from 3.42 to 41.88% while its impact is difficult to observe for GED scores and MSSIM due to the low number of compiled samples (i.e. low CSR) for zero-shot prompting. The effect of providing the schematic sheet size is negligible across all the metrics between Prompt 4 to Prompt 5. We select Prompt 2 for fine-tuning as it performs well out of all the zero-shot prompts and Prompt 4 due to the observed superiority in CSR. While, the effect of providing sheet size is not observable from this result, we also include Prompt 3 and 5 to examine how this affects the performance of the fine-tuned model. Furthermore, we used Llama-3.1-8B instruct model to leverage its knowledge base for text synthesis.
During fine-tuning, training and validation losses (Fig. 3) revealed overfitting (an increase in validation loss) after epoch 3. Since the cross entropy loss does not directly help quantify the correctness of the code and the generated schematics, we use GED $\overline{\text{GED score}}$ and $\overline{\text{MSSIM}}$ discussed in Section IV-D to evaluate the model performance across training epochs. GED $\overline{\text{GED score}}$s scaled by CSR for samples containing different numbers of components from epoch 0 (raw model of Llama-3.1-8B) to epoch 10 are shown in Fig. 4. We scale the scores by CSR for a fairer comparison since samples that are more difficult to generate (e.g. circuits with a large number of components) can become compilable during fine-tuning and be included in the calculation of the average score. This can prevent the scores from improving because these larger circuits tend to have lower scores than the smaller circuits as observed. The effect of different prompting techniques is not observable while the fine-tuning significantly improves the performance until the model overfits around epoch 3, which aligns with the observation from Fig. 3. Similarly, $\overline{\text{MSSIM}}$ scaled by CSR for different circuit size is shown in Fig. 5. In contrast to the previous observations, the scores improve until the last epoch. Furthermore, the impact of providing schematic sheet size is clearly reflected by the scores that are almost one-order of magnitude higher for (b) Prompt 3 and (d) Prompt 5 while one-shot prompting did not impact the performance.
We then choose Llama-3.1-8B fine-tuned with Prompt 3 using weights from epoch 6 as our best model and call it $Schemato$ since it demonstrated the best combination of scores in Fig. 4 and Fig. 5. We evaluate this model by comparing its performance with the Llama-3.1-8B baselines and GPT-4o, the latest and most powerful model from OpenAI. Table I shows that $Schemato$ outperforms the other models with 1.8×1.8$\times$1.8 × and 4.3×4.3$\times$4.3 × higher GED $\overline{\text{GED score}}$ and $\overline{\text{MSSIM}}$ scaled by CSR than the best performing pretrained models. Furthermore, $Schemato$ scores $76\%$ CSR compared to $63\%$ scored by the best performing pretrained model. Table II shows the ablation study of Llama-3.1-8B instruct fine-tuned with different prompts. There is no significant difference in the scores between the models fine-tuned with the zero-shot and one-shot prompts. In contrast, the dominance of the models fine-tuned with Prompt 3 ($Schemato$) and Prompt 5 in $\overline{\text{MSSIM}}$ and CSR showcases the effectiveness of including sheet size in the prompt.
Figure 4: GED $\overline{\text{GED score}}$s scaled by CSR evaluated with the test set. Llama-3.1-8B is fine-tuned with (a) Prompt 2, (b) Prompt 3, (c) Prompt 4, and (d) Prompt 5. The missing data points indicate that no generated samples are compilable (i.e. $\text{CSR}=0$). The number of components in the test circuits is denoted by $N_{c}$ and the number of reference samples composed of $N_{c}$ components is denoted by $N_{s}$.
Figure 5: $\overline{\text{MSSIM}}$ scaled by CSR evaluated with the test set. Llama-3.1-8B is fine-tuned with (a) Prompt 2, (b) Prompt 3, (c) Prompt 4, and (d) Prompt 5. The missing data points indicate that no generated samples are compilable (i.e. $\text{CSR}=0$). The number of components in the test circuits is denoted by $N_{c}$ and the number of reference samples composed of $N_{c}$ components is denoted by $N_{s}$.
@@
T1 N002 0 N003 0 Td=50n Z0=50
@@
(d) Llama 3.1 instruct
with Prompt 5
Figure 6: Netlist & LTSpice schematics of a transformer circuit
@@
(c) $Schemato$
(d) Llama 3.1 instruct
with Prompt 5
(e) GPT-4o
with Prompt 5
Figure 7: Netlist & LTSpice schematics of an RC band-pass filter
@@ VI Conclusions & Future work
In this paper, we introduced and evaluated $Schemato$, a fine-tuned LLM designed for converting circuit netlists into interpretable visual schematics. Our experimental results demonstrate that $Schemato$ significantly outperforms state-of-the-art LLMs, including Llama-3.1-8B and GPT-4o, across all the graph-based, image-based, compilation-based, and text-based metrics. These findings underscore the potential of $Schemato$ in reliably generating human-intuitive schematics from netlists, paving the way for enhanced interpretability and utility in ML-based analog circuit synthesis.
Our experiment in Fig. 4 show that for circuits with more than 5 components, $Schemato$ struggles to generate schematics with accurate connectivity. This is expected as there are more possible variations of circuits with different connectivity and components for these larger circuits, making it harder for $Schemato$ to generalize. The .asc file format further complicates generalization capabilities since it only specifies the components’ name and coordinate, but not symbol size. This means that $Schemato$ has to learn their symbol size from how wires are placed around the components across our training samples and therefore the model cannot generate schematics containing less familiar or unfamiliar components. The analysis on our dataset revealed that 5,227 out of the 9,907 training samples contain components that appear less than 10 times in our training set. Hence, more than half is composed of components of low familiarity. This highlights the need for a larger dataset with circuits composed of more generic and constrained set of components that are used in various configurations.
To address this challenge, future improvements to $Schemato$ will focus on enriching the dataset using advanced data augmentation techniques. For instance, we could combine two schematics into a new one using a technique similar to mixup used in vision [32]. To improve the generalizability of $Schemato$ to large and unseen circuit topologies, we will investigate the decomposition of larger circuits into smaller, more manageable subcircuits, thereby enhancing the ability of the model to learn hierarchical structures and the interplay between subcircuits. In addition, we plan to further expand the dataset by converting widely available schematic images into netlists and digitized schematics, using a method similar to that implemented in [33].